# {centering} Inflation models with Peccei-Quinn symmetry and axion kinetic misalignment

### v1

@@
{centering} Hyun Min Lee, Adriana G. Menkara, Myeong-Jung Seong,
and Jun-Ho Song
Department of Physics, Chung-Ang University, Seoul 06974, Korea.
We propose a consistent framework with the $U(1)$ Peccei-Quinn (PQ) symmetry for obtaining the initial condition for axion kinetic misalignment from inflation. We introduce a PQ complex scalar field and an extra Higgs doublet, which are conformally coupled to gravity, and three right-handed neutrinos for the seesaw mechanism. In the DFSZ type scenarios for the axion, we obtain the PQ anomalies from the Standard Model quarks carrying nonzero PQ charges in some of two Higgs doublet models, solving the strong CP problem by the QCD potential for the axion. Assuming that the PQ symmetry is explicitly violated in the scalar potential by quantum gravity effects, we show that a sufficiently large initial axion velocity can be obtained at the end inflation while avoiding the axion quality problem. As inflation is driven by the radial distance from the origin in the space of scalar fields close to the pole of the kinetic terms in the Einstein frame, we obtain successful inflationary predictions and set the initial axion velocity at the end of inflation. Focusing on the pure PQ inflation with a small running quartic coupling for the PQ field, we discuss the post-inflationary dynamics for the inflaton and the axion. As a result, we show that a sufficiently high reheating temperature, can be obtained dominantly from the Higgs-portal couplings to the PQ field, while being consistent with axion kinetic misalignment, the stability for the Higgs fields during inflation and the non-restoration of the PQ symmetry after reheating.
@@
Email: amenkara@cau.ac.kr
@@ 1 Introduction
The QCD axion is a pseudo-Goldstone boson of the $U(1)$ Peccei-Quinn (PQ) symmetry and it solves the strong CP problem in the Standard Model (SM) by the dynamical relaxation mechanism during the cosmological evolution after the QCD phase transition [1]. It is also a candidate for decaying cold dark matter, with feeble interactions to gluons and photons and potentially other SM particles [2, 3]. The relic abundance for the axion is determined by the misalignment mechanism where the axion is assumed to be displaced from the origin of the axion potential during or after inflation and becomes a coherently oscillating cold dark matter after the QCD phase transition.
@@ 2.2 Two Higgs doublet models
$\mathcal{L}^{(1)}_{\nu}=-Y_{\nu}\overline{l}_{L}\tilde{H}_{2}N_{R}-\frac{1}{2}\lambda_{N}\overline{N^{c}_{R}}\Phi N_{R}.$ (9)
@@
$\mathcal{L}^{(2)}_{\nu}=-Y_{\nu}\overline{l}_{L}\tilde{H}_{1}N_{R}-\frac{1}{2}\lambda_{N}\overline{N^{c}_{R}}\Phi N_{R}$ (10)
@@ 5.1 Inflaton condensate
When inflation is driven by the radial mode of the PQ field, the inflaton field value at the end of inflation is given by $\phi_{\rm end}=\sqrt{\frac{3}{8}}\ln\big{(}\frac{1}{6}(35+8\sqrt{19})\big{)}\simeq 1.50$ from ${\ddot{a}}=0$, for which the inflaton energy density at the end of inflation is given by $\rho_{\rm end}=\frac{3}{2}V_{E}(\phi_{\rm end})$. The post-inflationary potential of the inflaton for $|\Phi|\ll\sqrt{6}$ takes $V_{E}(\phi)\simeq\frac{1}{4}\lambda_{\Phi}\phi^{4}$, so the equation of state for the inflaton during reheating is radiative-like, i.e. $w_{\phi}=p_{\phi}/\rho_{\rho}=\frac{1}{3}$.
@@ 5.2 Reheating
${\cal L}_{\rm int}\supset-\sum_{i=1,2}\lambda_{i\Phi}|H_{i}|^{2}|\Phi|^{2}-2^{p/2-1}\kappa_{p}H^{\dagger}_{1}H_{2}\Phi^{p}-\frac{1}{2}\lambda_{N}\overline{N^{c}_{R}}\Phi N_{R}+{\rm h.c.},$ (88)
@@
$|M^{N}_{n}|^{2}=\lambda^{2}_{N}\phi^{2}_{0}|{\cal P}_{n}|^{2}E^{2}_{n}(\beta^{N}_{n})^{2},$ (105)
with $\beta^{N}_{n}=\sqrt{1-\frac{4m^{2}_{N_{R}}}{E^{2}_{n}}}$ and $m_{N}=\frac{1}{\sqrt{2}}\lambda_{N}\phi$ being the effective masses for the RH neutrinos. After averaging over oscillations, we get
$\langle\Gamma_{\phi\to N_{R}N_{R}}\rangle=\frac{9\lambda^{2}_{N}\omega^{3}}{8\pi m^{2}_{\phi}}\,\Sigma_{N}\Bigg{\langle}\bigg{(}1-\frac{4m^{2}_{N}}{\omega^{2}n^{2}}\bigg{)}^{3/2}\Bigg{\rangle},$ (106)
with $\Sigma_{N}=\sum_{n=1}^{\infty}n^{3}|{\cal P}_{n}|^{2}$. We also note that the Yukawa couplings of the RH neutrinos to the PQ field must be chosen to be small enough, namely, $\lambda_{N}\lesssim 10^{-3}$, in order to make the running effects on the quartic coupling $\lambda_{\Phi}$ ignorable. We note that as $4m^{2}_{N_{R}}/\omega^{2}\sim 2.79y^{2}_{N}/\lambda_{\Phi}$, the decay of the lowest inflaton modes into a pair of RH neutrinos are kinematically open if $|y_{N}|\lesssim 0.60\sqrt{\lambda_{\Phi}}\sim 2\times 10^{-6}$ for $\lambda_{\Phi}=1.1\times 10^{-11}$.
@@
$|{M}^{N^{\prime}}_{n}|^{2}=\frac{32\lambda^{4}_{N}m^{2}_{N}\phi^{4}_{0}({\beta}^{N^{\prime}}_{n})^{2}}{E^{2}_{n}}$ (108)
@@
$\langle\Gamma_{\phi\phi\to N_{R}N_{R}}\rangle=\frac{36\lambda^{4}_{N}\phi^{2}_{0}}{\pi m^{2}_{\phi}\omega}\,{\hat{\Sigma}}_{N}\bigg{\langle}m^{2}_{N}\bigg{(}1-\frac{m^{2}_{N}}{n^{2}\omega^{2}}\bigg{)}^{3/2}\bigg{\rangle},$ (109)
@@ 5.3 Reheating temperature
$T_{\rm RH}|_{\rm decay}$ $\simeq$ $5.2\times 10^{5}\,{\rm GeV}\bigg{(}\frac{100}{g_{*}(T_{\rm reh})}\bigg{)}^{1/4}\bigg{(}\frac{{\rm max}[y_{N},\kappa_{1}/\omega]}{10^{-4}}\bigg{)}^{2}\bigg{(}\frac{\lambda_{\Phi}}{10^{-11}}\bigg{)}^{1/4},$ (114)
$T_{\rm RH}|_{\rm scattering}$ $\simeq$ $3.0\times 10^{11}\,{\rm GeV}\bigg{(}\frac{100}{g_{*}(T_{\rm reh})}\bigg{)}^{1/4}\bigg{(}\frac{{\rm max}[\lambda_{H_{i}\Phi},\kappa_{2},\sqrt{2}y^{2}_{N}m_{N}/\omega]}{10^{-7}}\bigg{)}^{2}\bigg{(}\frac{10^{-11}}{\lambda_{\Phi}}\bigg{)}^{3/4}\,.$ (115)
@@ 6.2 Axion relic density
the pressure for the domain walls becomes dominant over the energy density of the domain walls before Big Bang Nucleosynthesis, as far as $\Delta V\gtrsim\frac{\sigma^{2}}{M^{2}_{P}}(t_{*}/0.1\,{\rm s})$ where $\sigma$ is the tension of domain walls and $t_{*}$ is the time at which the energy density of the domain walls dominates the radiation energy density, given by $t_{*}\sim\frac{M^{2}_{P}}{\sigma^{3}}$ [22]. Thus, for $\sigma\sim\Lambda^{3}_{\rm QCD}$, there is no domain wall problem as far as $c\gtrsim 10^{-13}$, which is satisfied in a consistent parameter space for $l$ and $f_{a}$ where the condition for the axion quality in eq. (42) and the CMB bound in eq. (85) are satisfied simultaneously. For instance, it is sufficient to take $l=10(13)$ for $f_{a}=10^{9}(10^{12})\,{\rm GeV}$. Otherwise, in the region of the parameter space with $T_{\rm reh}>T_{\rm restore}$ in Fig. 1, it is relevant to consider the production of axion domain walls and their impact on the axion relic density.
@@ Acknowledgments
We are supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567).

### v2

@@
{centering} Hyun Min Lee1†, Adriana G. Menkara1,2♯, Myeong-Jung Seong1‡,
and Jun-Ho Song1⋆
1Department of Physics, Chung-Ang University, Seoul 06974, Korea.
2Deutsches Elektronen-Synchrotron DESY, Notkestrasse 85, 22607 Hamburg, Germany.
We propose a consistent framework with the $U(1)$ Peccei-Quinn (PQ) symmetry for obtaining the initial condition for axion kinetic misalignment from inflation. We introduce a PQ complex scalar field and an extra Higgs doublet, which are conformally coupled to gravity, and three right-handed neutrinos for the seesaw mechanism. In the DFSZ type scenarios for the axion, we obtain the PQ anomalies from the Standard Model quarks carrying nonzero PQ charges in some of two Higgs doublet models, solving the strong CP problem by the QCD potential for the axion. Assuming that the PQ symmetry is explicitly violated in the scalar potential by quantum gravity effects, we show that a sufficiently large initial axion velocity can be obtained before the QCD phase transition while avoiding the axion quality problem. As inflation is driven by the radial distance from the origin in the space of scalar fields close to the pole of the kinetic terms in the Einstein frame, we obtain successful inflationary predictions and set the initial axion velocity at the end of inflation. Focusing on the pure PQ inflation with a small running quartic coupling for the PQ field, we discuss the post-inflationary dynamics for the inflaton and the axion. As a result, we show that a sufficiently high reheating temperature can be obtained dominantly from the Higgs-portal couplings to the PQ field, while being consistent with axion kinetic misalignment, the stability for the Higgs fields during inflation and the non-restoration of the PQ symmetry after reheating.
@@
Email: adriana.menkara@desy.de
@@ 1 Introduction
The QCD axion is a pseudo-Goldstone boson of the $U(1)$ Peccei-Quinn (PQ) symmetry and it solves the strong CP problem in the Standard Model (SM) by the dynamical relaxation mechanism during the cosmological evolution after the QCD phase transition [1]. It is also a candidate for decaying cold dark matter [2], with feeble interactions to gluons and photons and potentially other SM particles [3, 4]. The relic abundance for the axion is determined by the misalignment mechanism [2] where the axion is assumed to be displaced from the origin of the axion potential during or after inflation and becomes a coherently oscillating cold dark matter after the QCD phase transition.
@@ 2.2 Two Higgs doublet models
$\mathcal{L}^{(1)}_{\nu}=-Y_{\nu}\overline{l}_{L}\tilde{H}_{2}N_{R}-\frac{1}{2}y_{N}\overline{N^{c}_{R}}\Phi N_{R}.$ (9)
@@
$\mathcal{L}^{(2)}_{\nu}=-Y_{\nu}\overline{l}_{L}\tilde{H}_{1}N_{R}-\frac{1}{2}y_{N}\overline{N^{c}_{R}}\Phi N_{R}$ (10)
@@ 4.3 Slow-roll inflation and axion rotation
We remark on the multiple dynamics other than the radial component of the PQ field. In both the PQ-Higgs mixed inflation and the pure PQ inflation, it was shown that the field directions orthogonal to the inflaton field can receive masses much greater than the Hubble scale during inflation. But, if the masses of the non-inflaton fields are comparable to or smaller than the Hubble scale in a certain parameter space, we need to consider the dynamics of multiple fields during inflation, for instance, for the non-Gaussianity of CMB [15]. However, the detailed discussion on the dynamics of multiple fields and the non-Gaussianity is beyond the scope of our work, so we postpone it in a future work.
@@ 5.1 Inflaton condensate
When inflation is driven by the radial mode of the PQ field, the inflaton field value at the end of inflation is given by $\phi_{\rm end}=\sqrt{\frac{3}{8}}\ln\big{(}\frac{1}{6}(35+8\sqrt{19})\big{)}\simeq 1.50$ from ${\ddot{a}}=0$, for which the inflaton energy density at the end of inflation is given by $\rho_{\rm end}=\frac{3}{2}V_{E}(\phi_{\rm end})$. The post-inflationary potential of the inflaton for $|\Phi|\ll\sqrt{3}$ takes $V_{E}(\phi)\simeq\frac{1}{4}\lambda_{\Phi}\phi^{4}$, so the equation of state for the inflaton during reheating is radiative-like, i.e. $w_{\phi}=p_{\phi}/\rho_{\rho}=\frac{1}{3}$.
@@ 5.2 Reheating
${\cal L}_{\rm int}\supset-\sum_{i=1,2}\lambda_{i\Phi}|H_{i}|^{2}|\Phi|^{2}-2^{p/2-1}\kappa_{p}H^{\dagger}_{1}H_{2}\Phi^{p}-\frac{1}{2}y_{N}\overline{N^{c}_{R}}\Phi N_{R}+{\rm h.c.},$ (88)
@@
$|M^{N}_{n}|^{2}=y^{2}_{N}\phi^{2}_{0}|{\cal P}_{n}|^{2}E^{2}_{n}(\beta^{N}_{n})^{2},$ (105)
with $\beta^{N}_{n}=\sqrt{1-\frac{4m^{2}_{N}}{E^{2}_{n}}}$ and $m_{N}=\frac{1}{\sqrt{2}}y_{N}\phi$ being the effective masses for the RH neutrinos. After averaging over oscillations, we get
$\langle\Gamma_{\phi\to N_{R}N_{R}}\rangle=\frac{9y^{2}_{N}\omega^{3}}{8\pi m^{2}_{\phi}}\,\Sigma_{N}\Bigg{\langle}\bigg{(}1-\frac{4m^{2}_{N}}{\omega^{2}n^{2}}\bigg{)}^{3/2}\Bigg{\rangle},$ (106)
with $\Sigma_{N}=\sum_{n=1}^{\infty}n^{3}|{\cal P}_{n}|^{2}$. We also note that the Yukawa couplings of the RH neutrinos to the PQ field must be chosen to be small enough, namely, $y_{N}\lesssim 10^{-3}$, in order to make the running effects on the quartic coupling $\lambda_{\Phi}$ ignorable. We note that as $4m^{2}_{N}/\omega^{2}\sim 2.79y^{2}_{N}/\lambda_{\Phi}$, the decay of the lowest inflaton modes into a pair of RH neutrinos are kinematically open if $|y_{N}|\lesssim 0.60\sqrt{\lambda_{\Phi}}\sim 2\times 10^{-6}$ for $\lambda_{\Phi}=1.1\times 10^{-11}$.
@@
$|{M}^{N^{\prime}}_{n}|^{2}=\frac{32y^{4}_{N}m^{2}_{N}\phi^{4}_{0}({\beta}^{N^{\prime}}_{n})^{2}}{E^{2}_{n}}$ (108)
@@
$\langle\Gamma_{\phi\phi\to N_{R}N_{R}}\rangle=\frac{36y^{4}_{N}\phi^{2}_{0}}{\pi m^{2}_{\phi}\omega}\,{\hat{\Sigma}}_{N}\bigg{\langle}m^{2}_{N}\bigg{(}1-\frac{m^{2}_{N}}{n^{2}\omega^{2}}\bigg{)}^{3/2}\bigg{\rangle},$ (109)
@@ 5.3 Reheating temperature
$T_{\rm RH}|_{\rm decay}$ $\simeq 5.2\times 10^{5}\,{\rm GeV}\bigg{(}\frac{100}{g_{*}(T_{\rm reh})}\bigg{)}^{1/4}\bigg{(}\frac{{\rm max}[y_{N},\kappa_{1}/\omega]}{10^{-4}}\bigg{)}^{2}\bigg{(}\frac{\lambda_{\Phi}}{10^{-11}}\bigg{)}^{1/4},$ (114)
$T_{\rm RH}|_{\rm scattering}$ $\simeq 3.0\times 10^{11}\,{\rm GeV}\bigg{(}\frac{100}{g_{*}(T_{\rm reh})}\bigg{)}^{1/4}\bigg{(}\frac{{\rm max}[\lambda_{H_{i}\Phi},\kappa_{2},\sqrt{2}y^{2}_{N}m_{N}/\omega]}{10^{-7}}\bigg{)}^{2}\bigg{(}\frac{10^{-11}}{\lambda_{\Phi}}\bigg{)}^{3/4}\,.$ (115)
@@ 6.2 Axion relic density
the pressure for the domain walls becomes dominant over the energy density of the domain walls before Big Bang Nucleosynthesis, as far as $\Delta V\gtrsim\frac{\sigma^{2}}{M^{2}_{P}}(t_{*}/0.1\,{\rm s})$ where $\sigma$ is the tension of domain walls and $t_{*}$ is the time at which the energy density of the domain walls dominates the radiation energy density, given by $t_{*}\sim\frac{M^{2}_{P}}{\sigma}$ [24]. Thus, for $\sigma\sim\Lambda^{2}_{\rm QCD}f_{a}$, there is no domain wall problem as far as $c\gtrsim 10^{-3}(f_{a}/10^{9}\,{\rm GeV})$, which is satisfied in a consistent parameter space for $l$ and $f_{a}$ where the condition for the axion quality in eq. (42) and the CMB bound in eq. (85) are satisfied simultaneously. For instance, it is sufficient to take $l=9(12)$ for $f_{a}=10^{9}(10^{12})\,{\rm GeV}$. Otherwise, in the region of the parameter space with $T_{\rm reh}>T_{\rm restore}$ in Fig. 1, it is relevant to consider the production of axion domain walls and their impact on the axion relic density.
@@
Before closing this section, we remark on the isocurvature perturbations of the axion generated during inflation [25]. The total axion relic density is composed of $Y_{a}=Y_{a,{\rm kin}}+Y_{a,{\rm mis}}$, with $sY_{a,{\rm kin}}=2m_{a}n_{\theta}$ and $sY_{a,{\rm mis}}=\frac{1}{2}m^{2}_{a}f^{2}_{a}\theta^{2}_{*}$. Then, the power spectrum of the isocurvature perturbation at the horizon exit of the mode $k_{*}$ during inflation depends only on $Y_{a,{\rm mis}}$ is given by
$P_{\rm iso}(k_{*})$ $=$ $\bigg{(}\frac{1}{Y_{a}}\frac{\partial Y_{a}}{\partial\theta_{*}}\bigg{)}^{2}\langle\delta\theta_{*}^{2}\rangle$ (132)
$=$ $\bigg{[}\frac{4}{\theta^{2}_{*}}\bigg{(}\frac{Y_{a,{\rm mis}}}{Y_{a}}\bigg{)}^{2}+\frac{1}{4}\bigg{(}\frac{1}{\epsilon_{\theta,*}}\frac{\partial\epsilon_{\theta,*}}{\partial\theta_{*}}\bigg{)}^{2}\bigg{(}\frac{Y_{a,{\rm kin}}}{Y_{a}}\bigg{)}^{2}\bigg{]}\langle\delta\theta_{*}^{2}\rangle$
where the fluctuation of the initial angle is given by
$\langle\delta\theta_{*}^{2}\rangle=\frac{1}{f^{2}_{a,{\rm eff}}}\,\bigg{(}\frac{H_{I}}{2\pi}\bigg{)}^{2}.$ (133)
with $f_{a,{\rm eff}}\equiv\sqrt{6}\big{|}\sinh\big{(}\frac{\phi_{*}}{\sqrt{6}M_{P}}\big{)}\big{|}$. Here, we took into account the non-canonical kinetic term for the fluctuation of $\theta$ in eq. (63) in the Einstein frame, so the effective axion decay constant becomes very large for $\phi_{*}\gg M_{P}$ during inflation. The fluctuation of $Y_{a,{\rm kin}}$ with respect to $\theta$ is nonzero, because the initial axion velocity during inflation depends on the initial value of $\theta$ through the explicit PQ breaking term in eq. (73), namely, the slow-roll parameter $\epsilon_{\theta}$ for $\theta$. Although the fluctuation of $Y_{a,{\rm kin}}$ also depends on the inflaton field $\phi$ through the PQ charge, $n_{\theta}$, in eq. (77), it does not contribute to the isocurvature perturbations because the inflaton fluctuation is adiabatic. The bound on the isocurvature perturbations from Planck satellite [12] is given by
$\beta_{\rm iso}\equiv\frac{P_{\rm iso}(k_{*})}{P_{\zeta}(k_{*})+P_{\rm iso}(k_{*})}<0.038,$ (134)
at $95\%$ C.L., with $P_{\zeta}(k_{*})=A_{s}=2.1\times 10^{-9}$ at $k_{*}=0.05\,{\rm Mpc}^{-1}$. Then, taking $Y_{a}\simeq Y_{a,{\rm kin}}$ and using
$\bigg{(}\frac{1}{\epsilon_{\theta,*}}\frac{\partial\epsilon_{\theta,*}}{\partial\theta}\bigg{)}^{2}=4(l-2k)^{2}\cot^{2}\Big{(}(l-2k)\theta_{*}+A_{0,l,k}\Big{)},$ (135)
we get
$\frac{4}{\theta^{2}_{*}}\bigg{(}\frac{Y_{a,{\rm mis}}}{Y_{a,{\rm kin}}}\bigg{)}^{2}+(l-2k)^{2}\cot^{2}\Big{(}(l-2k)\theta_{*}+A_{0,l,k}\Big{)}\lesssim 10.$ (136)
Here, we assumed a single explicit PQ breaking term at the order of $f^{l}_{a}$ in eq. (65) and used $\epsilon_{*}\simeq 2\times 10^{-4}$ for $N=60$ in eq. (82), for which $\phi_{*}\simeq 7.1M_{P}$ and the effective axion decay constant during inflation becomes $f_{a,{\rm eff}}\simeq 22M_{P}$. Then, the above result leads to $Y_{a,{\rm mis}}/Y_{a,{\rm kin}}\lesssim 5$ and $|(l-2k)\cot((l-2k)\pi+A_{0,l,k})|\lesssim 3$ for $\theta_{*}=\pi$, which are consistent with the dominance of the axion kinetic misalignment, as far as $|l-2k|\lesssim 3$ is taken or a mild fine-tuning between $\theta_{*}$ and the phase shift $A_{0,l,k}$ is made for a small value of $|\cot((l-2k)\theta_{*}+A_{0,l,k})|$ with a large $|l-2k|$. As a result, in our model, the isocurvature perturbations are sufficiently small due to a large value of the effective axion decay constant during inflation.
@@ Acknowledgments
We are supported in part by Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567). This research was supported by the Chung-Ang University research grant in 2019. AM acknowledges support by the DFG Emmy Noether Grant No. PI 1933/1-1.